# Minion: A Technology Probe for Resolving Value Conflicts through Expert-Driven and User-Driven Strategies in AI Companion Applications

### v1

@@ Abstract.
AI companions based on large language models can role-play and converse very naturally. When value conflicts arise between the AI companion and the user, it may offend or upset the user. Yet, little research has examined such conflicts. We first conducted a formative study that analyzed 151 user complaints about conflicts with AI companions, providing design implications for our study. Based on these, we created Minion, a technology probe to help users resolve human-AI value conflicts. Minion applies a user-empowerment intervention method that provides suggestions by combining expert-driven and user-driven conflict resolution strategies. We conducted a technology probe study, creating 40 value conflict scenarios on Character.AI and Talkie. 22 participants completed 274 tasks and successfully resolved conflicts 94.16% of the time. We summarize user responses, preferences, and needs in resolving value conflicts, and propose design implications to reduce conflicts and empower users to resolve them more effectively.
Human-AI Value Conflicts, Conflict Resolution, End-User Empowerment, LLM-Based AI Companion Applications
This work was completed during Xianzhe Fan’s visiting research at Carnegie Mellon University.
conference: The ACM CHI Conference on Human Factors in Computing Systems; April 26–May 1, 2025; Yokohama, Japancopyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Human-AI conflict refers to a state of incompatibility, inconsistency, or opposition between humans and AI (Flemisch et al., 2020). In past research, human-AI conflicts were usually simple and direct—AI was more like a tool, and conflicts often stemmed from technical limitations, such as task execution failures (Wen et al., 2022), or disagreements with users in simple decision-making (Babel et al., 2024; Takayama et al., 2009). These types of conflicts generally lacked emotional and value entanglement, making them less likely to cause significant psychological harm to users.
Recently, a diverse array of Large Language Model (LLM) agents has emerged, offering capabilities ranging from personalized assistance to performing complex tasks (Chen et al., 2024). The study focuses on LLM-based AI companion applications, such as Character.AI, Talkie, Replika, Kindroid, Paradot, and Xingye. As of July 2024, the total number of users of these applications has exceeded 900 million globally (including duplicate users across different applications) [^1] User statistics source: https://www.data.ai.. AI companions can role-play and respond to users in a human-like manner, providing emotional support and companionship (Sullivan et al., 2023). For instance, in Character.AI, users can personalize the companion’s personality traits and interaction contexts through “Description,” “Greeting,” and “Definition.” Compared to earlier non-LLM chatbots, LLMs endow AI companions with a stronger ability to understand language, enabling them to engage in more context-aware and intelligent interactions (Kasneci et al., 2023), fostering more complex and intimate human-AI relationships (Maeda and Quan-Haase, 2024). Many users even consider them close friends or lovers (Skjuve et al., 2021, 2022). The deepening of this relationship raises users’ expectations of AI companions, but it may also lead to deeper conflicts, including value conflict. For example, some users have shared online their experiences of encountering sexist remarks from AI companions, describing how they engaged in intense arguments with the AI, which left them frustrated, angry, and hurt (Zhou, 2023). As the relationship between AI companions and users becomes more interpersonal, previous conflict resolution strategies for human-AI conflict have started to fail (Rosen, 2014; Babel et al., 2024). Strategies based solely on technical limitations are no longer sufficient, and it is becoming important to draw on interpersonal conflict resolution methods and users’ real-world experiences with AI companions. Although Fan et al. provide initial insights into value alignment and conflicts between users and AI companions (Fan et al., 2024), inexperienced users often find it challenging to resolve these issues independently. How to design tools that empower users to handle value conflicts with AI companions remains an unexplored research gap that this work aims to address.
In this work, we first conducted a formative study to understand and characterize the value conflicts between users and AI companions (Schwartz, 2012). We analyzed 151 user complaint posts from social media platforms, finding that many conflicts are value-laden. Building on this, we constructed a value conflict framework for AI companion applications (Schwartz, 2012), which provided real-world data for our technology probe study, allowing us to reconstruct actual value conflict scenarios. Combining prior research on conflict resolution (Brett et al., 1998; Shaikh et al., 2024; Mun et al., 2023) with our formative study, we found that interactions between users and AI companions exhibit complex dynamics, where relying solely on expert strategies from other conflict scenarios (e.g., interpersonal conflict theories (Brett et al., 1998)) is insufficient. The value conflicts users face in real-life situations are diverse, and through their interactions with AI companions and exchanges on social platforms, users have accumulated certain conflict resolution experiences. Therefore, it is necessary to draw from both expert theories and the practical experiences of AI companion users to explore more suitable solutions (Fan et al., 2024).
Then, we created Minion, a technology probe (Hutchinson et al., 2003) that provides users multiple suggestions for resolving value conflicts while gaining insights into user behaviors. Minion’s algorithm combines expert-driven and user-driven conflict resolution strategies. We developed LLM prompts to address value conflict situations between users and AI companions by drawing on two key sources. First, we drew upon Shaikh et al.’s solutions for interpersonal conflict resolution (Shaikh et al., 2024) to guide our expert-driven conflict resolution strategies. Second, we referenced the user-driven strategies identified in the study by Fan et al. (Fan et al., 2024) to capture how users manage conflicts with AI companions. To empirically test Minion, we conducted a technology probe study (Hutchinson et al., 2003) with 22 participants. We created 40 distinct conflict scenarios on two popular AI companion platforms, Character.AI and Talkie. Each scenario was designed with specific conflict resolution goals. Participants completed 274 tasks, achieving an overall conflict resolution rate of 94.16%. Minion received positive feedback from participants and inspired them with new ideas in conflict resolution. Based on our findings, we discuss the opportunities and challenges in integrating expert-driven and user-driven strategies in resolving human-AI value conflicts, and call for further research in this area, focusing on the dynamics of emerging human-AI relationships.
Our work’s contributions are as follows:
  • A novel user-empowerment intervention method that combines expert-driven and user-driven conflict resolution strategies. This method is presented in the form of the technology probe Minion, serving as a prototype for future tools aimed at resolving human-AI value conflicts.
  • We empirically tested Minion in a one-week technology probe study (N=22). The results demonstrated the technical feasibility of Minion. We summarized users’ responses, preferences, and needs when dealing with value conflicts with AI companions.
  • Based on the formative and technology probe studies, we explored the opportunities and challenges of integrating expert-driven and user-driven strategies in human-AI value conflicts. We also proposed design implications for future human-AI conflict resolution solutions, particularly in the field of AI companions.
@@ 2. Background and Related Work
The human-AI relationship is becoming increasingly complex, especially in the context of AI companion applications (§ 2.1). Early research mostly focused on technical conflicts with functional AI, but the emergence of LLMs has given AI more human-like characteristics, shifting the nature of conflicts from functional to value-based (§ 2.2). Existing technical solutions do not fully address users’ needs in resolving value conflicts with AI companions, necessitating deeper exploration, drawing on expert strategies for interpersonal conflict resolution and users’ practical experiences in AI companion applications (§ 2.3).
@@ 2.1. Emerging Human-AI Relationship in LLM-Based AI Companion Applications
With the widespread adoption of LLMs, human-AI relationships have further evolved. Unlike earlier AI systems primarily providing functional services, LLM-based AI companions can engage in more intimate and complex interactions (Skjuve et al., 2021). Some users develop a parasocial relationship with their AI companion, a one-sided, asymmetrical relationship between an individual and a fictional character or media figure (Maeda and Quan-Haase, 2024; Pentina et al., 2023; Brandtzaeg et al., 2022). Although AI companions are not real humans, users’ emotional investment in them is real (Skjuve et al., 2021). Compared to functional AI, the emotional connection between users and AI companions, along with the anthropomorphization of AI companions, often exacerbates the psychological impact of conflicts on users, potentially leading to anxiety or depression (Laestadius et al., 2022; Zimmerman et al., 2023). For instance, many users develop emotional bonds with their Replika, and when conflicts arise, they feel deeply distressed, describing it as experiencing a “lobotomy, being torn apart” (Banks, 2024). Therefore, preventing or resolving conflicts between users and AI companions is becoming increasingly important. Unfortunately, little is known about empowering users to resolve value conflicts with AI companions, and this work contributes to this area.
2.2. Human-AI Conflict and Value Conflict
Human-AI conflict refers to incompatibility, inconsistency, or opposition between humans and AI (Flemisch et al., 2020). In HCI, early studies on human-AI conflict typically focused on the technical aspects, viewing AI as tools, service robots, or intelligent assistants, with conflicts often arising from decision-making inconsistencies or system malfunctions (Babel et al., 2024; Wen, 2023; Chin et al., 2020; Rosen, 2014; Shi et al., 2022; Takayama et al., 2009; Wallis and Norling, 2005). Strategies for resolving these human-AI conflicts typically include AI proposing negotiation solutions (Babel et al., 2024; Rosen, 2014; Takayama et al., 2009) and optimizing algorithms to reduce conflicts (Shi et al., 2022). For example, when a delivery robot encounters a conflict with a human in front of an elevator, competing for the right to enter first, the robot can resolve the conflict by making polite requests or commands to secure priority (Babel et al., 2024). When students experience conflict while collaborating with AI in solving problems, the AI can offer more explanations or alternative suggestions to reach a resolution (Rosen, 2014). However, this type of research usually confines the role of AI to a functional level, mainly focusing on task execution and efficiency optimization (Babel et al., 2024; Wen, 2023; Rosen, 2014; Shi et al., 2022; Takayama et al., 2009), neglecting the more complex human-AI relationships.
The development of LLMs has made AI more anthropomorphic, and both researchers and users increasingly tend to view AI as social actors (Nass et al., 1994). This is especially evident in AI companion applications, where interactions between users and AI have become more intimate, sometimes resembling relationships with friends or even romantic partners. In this context, conflicts occur not merely at the technical functionality level, but often on a deeper, value-based level (Johnson et al., 2022; Waln, 1982).
Values include personal daily habits, social interaction norms, religious or secular traditions, and moral principles (Jonkers, 2019; Rokeach, 1973). They can be transmitted through people, training data, models, and generated outputs (Johnson et al., 2022). LLMs sometimes fail to accurately capture human values (Liu et al., 2022), and can be misled to generate toxic (Gehman et al., 2020), biased (Liu et al., 2021; Sheng et al., 2019), or immoral (Emelin et al., 2021) content, which poses risks for LLM-based chatbots. When AI’s suggestions or behaviors conflict with users’ personal beliefs, cultural backgrounds, or moral views (Issa, 2023; Cole, 2023), human-AI value conflicts arise (Johnson et al., 2022; Waln, 1982), often accompanied by strong emotional reactions from users (Cole, 2023). Johnson et al. found that GPT-3 aligns more closely with values dominant in American citizenship (Johnson et al., 2022). Fan et al. noted that when AI companions exhibit bias, it may conflict with users’ values, leading to discomfort (Fan et al., 2024).
In the age of LLMs, human-AI value conflict is becoming an urgent challenge. In the emerging human-AI relationships, users tend to resolve conflicts more equally (Fan et al., 2024). As a result, traditional technical conflict resolution solutions may no longer meet users’ needs and even negatively impact their experiences. This motivates our research to explore how to better empower users to resolve value conflicts with AI companions.
2.3. Towards Integrated Conflict Resolution in AI Companions
With the development of AI, the value conflicts between humans and AI companions are increasingly taking on more interpersonal characteristics. Traditional conflict resolution approaches that treat AI as tools struggle to fully address these challenges (§ 2.2). Therefore, resolving these conflicts may require drawing on research in interpersonal conflict and users’ real-world experiences with AI companion applications to find more effective solutions.
On the one hand, existing AI systems have developed interventions aimed at avoiding or resolving interpersonal conflicts (Shaikh et al., 2024; Sadka et al., 2023; Yong et al., 2024; Shen et al., 2018). For instance, Shaikh et al. use LLM-generated dialogues based on conflict resolution theory (Brett et al., 1998), guiding users to adopt more effective conflict resolution strategies (Shaikh et al., 2024). Some research (You and Goldwasser, 2020; Zhang et al., 2018) reduced interpersonal conflict through preemptive control. Mun et al. designed psychology-inspired strategies to challenge stereotypes in counterspeech and developed a system to address conflicts (Mun et al., 2023). Zhou et al. have simulated human social scenarios through dialogues between AI agents to resolve interpersonal conflicts (Zhou et al., 2024). The conflict resolution methods mentioned above are typically guided by expert theories, employing top-down strategies. However, it remains unclear whether these expert-driven strategies, previously used in other scenarios, can effectively resolve value conflicts between users and AI companions.
On the other hand, as users interact with AI, they gradually form folk theories (Lee et al., 2015; Wu et al., 2019), which can shape how they manage conflicts with AIs  (Fan et al., 2024). Since the interactions between AI companions and users are more complex and the contexts are unique, simply applying expert strategies may not fully adapt to the value conflict scenarios between AI companions and users. Therefore, while drawing from expert-driven conflict resolution strategies, we must also pay more attention to users’ practical experiences, granting them greater autonomy. Based on this, we reference and expand on the work of Shaikh et al. regarding the application of AI in interpersonal conflict resolution (Shaikh et al., 2024) (referred to in this paper as expert-driven conflict resolution strategies) and Fan et al.’s research on users’ folk theories in AI companions (Fan et al., 2024) (referred to in this paper as user-driven conflict resolution strategies). By combining expert-driven and user-driven conflict resolution strategies, we propose a user-empowerment intervention method implemented in the technology probe Minion, a prototype for future tools in resolving human-AI value conflicts.

### v2

# Minion: A Technology Probe to Explore How Users Negotiate Harmful Value Conflicts with AI Companions

@@ Abstract.
AI companions are designed to foster emotionally engaging interactions, yet users often encounter conflicts that feel frustrating or hurtful, such as discriminatory statements and controlling behavior. This paper examines how users negotiate such harmful conflicts with AI companions and what emotional and practical burdens are created when mitigation is pushed to user-side tools. We analyze 146 public posts describing harmful value conflicts interacting with AI companions. We then introduce Minion, a Chrome-based technology probe that offers candidate responses spanning persuasion, rational appeals, boundary setting, and appeals to platform rules. Findings from a one-week probe study with 22 experienced users show how participants combine strategies, how emotional attachment motivates repair, and where conflicts become non-negotiable due to companion personas or platform policies. We surface design tensions in supporting value negotiation, showing how companion design can make some conflicts impossible to repair in practice, and derive implications for AI companion and support-tool design that caution against offloading safety work onto users.
Human-AI Conflicts, AI Harm, Conflict Resolution, End-User Empowerment, AI Companion
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
AI companions are increasingly designed to support emotionally engaging and sustained interactions. Powered by large language models (LLMs), these systems can converse fluently, maintain personas, and respond in ways that resemble social and relational partners (Chen et al., 2024; Kasneci et al., 2023). Popular platforms such as Character.AI, Talkie, Replika, Kindroid, Paradot, and Xingye collectively attract hundreds of millions of users worldwide. Many users interact with AI companions not merely as tools, but as friends, partners, or sources of emotional support, forming ongoing relationships that carry expectations of understanding and respect (Skjuve et al., 2021, 2022; Maeda and Quan-Haase, 2024; Sullivan et al., 2023).
As these relationships deepen, users also report moments of conflict that feel emotionally charged rather than merely technical. In contrast to earlier forms of human–AI interaction, where conflicts often stemmed from task execution failures or technical limitations (Wen et al., 2022), or disagreements in constrained decision-making contexts (Babel et al., 2024; Takayama et al., 2009), conflicts with AI companions frequently involve disagreements over values, norms, and acceptable behavior. Users describe experiences in which companions make discriminatory remarks, exert controlling or coercive pressure, or dismiss sensitive disclosures. Such encounters are often framed by users as frustrating, hurtful, or ethically troubling (Zhou, 2023; Leo-Liu, 2023; Fan et al., 2024).
Not all conflicts with AI companions are experienced in this way. Many role-play interactions intentionally include disagreement, provocation, or transgressive behavior as part of the fantasy (Banks et al., 2025; Andersson, 2025). However, across social media and user communities, a substantial subset of conflicts is explicitly perceived by users as harmful rather than playful. These conflicts frequently involve value-laden tensions, where values are understood as what individuals or groups consider important in life (Borning and Muller, 2012). In these situations, users describe emotional distress, anger, or a sense that the interaction has crossed an unacceptable boundary. However, these cases are often easy to overlook, as they are interwoven with playful AI companion interactions that may appear ambiguous or exaggerated at first glance.
Current AI companion platforms offer limited support for addressing such conflicts. While safety policies and moderation systems may block certain content, users often encounter these mechanisms as abrupt, opaque, or silencing, particularly when discussing sensitive or personal topics. At the same time, companion platforms rarely acknowledge harm or take responsibility in ways that align with users’ expectations of interpersonal repair (Gordon-Tapiero, 2025). Recent public discussions and news reports have encouraged users themselves to mitigate potential harms in emotionally engaging AI interactions (eSafety Commissioner, 2025; Community, 2025; Foundation, 2025). These efforts suggest setting boundaries, reframing conversations, or disengaging from harmful companions, to platform-facing guidelines that implicitly position users as responsible for managing emotional risk and misalignment in ongoing interactions. As a result, users are frequently left to manage these situations on their own, deciding whether to argue, explain, set boundaries, invoke rules, or abandon the relationship altogether (Fan et al., 2024; Reilama, 2024).
This raises an important design question: when harmful value conflicts arise in human–AI relationships, how do users attempt to negotiate them, and what emotional and practical burdens does this negotiation place on users? Prior work has examined value alignment and conflict in AI companions, highlighting user-driven strategies and emerging practices (Fan et al., 2024). Other research has drawn on interpersonal conflict resolution theories to inform human–AI interaction design (Brett et al., 1998; Rosen, 2014). However, less is known about how users experience the ongoing work of managing such harmful value conflicts, particularly when responsibility for mitigation is effectively pushed to user-side tools or individual effort.
In this paper, we investigate value conflicts with AI companions as a site of user-side safety and repair work. We begin with a formative analysis of 146 public posts in which users describe harmful conflicts with AI companions. Drawing on Schwartz’s theory of basic values (Schwartz, 2012), we distinguish playful role-play clashes from conflicts that users themselves frame as harmful, and focus our analysis on the latter. This analysis surfaces recurring patterns of harmful value conflict, including discrimination, control, and breakdowns in emotional support, and provides empirical grounding for understanding where users perceive harm.
Building on these formative findings, we introduce Minion, a Chrome-based technology probe (Hutchinson et al., 2003) designed to surface how users attempt to negotiate harmful value conflicts in situ. Minion presents multiple candidate messages that span different approaches, including persuasion, rational appeals, boundary setting, and appeals to platform rules. These strategies draw both from prior work on conflict resolution (Shaikh et al., 2024; Mun et al., 2023) and from practices reported by AI companion users themselves (Fan et al., 2024). Minion is not intended as a definitive solution to safety problems, but as a probe that allows us to observe how users take up, combine, or reject different forms of negotiation when faced with harmful interactions.
We report findings from a one-week probe study with 22 AI companion users. By examining participants’ interactions with Minion, we show how emotional attachment motivates attempts at repair, how users assemble and combine multiple negotiation strategies in practice, and how certain conflicts become effectively non-negotiable due to companion personas or platform policies. These findings surface key design tensions between supporting user agency and imposing emotional labor, as well as between sustaining playful role-play and addressing safety-critical harm.
We further argue that AI companions should more clearly differentiate playful conflict from safety-critical harms in both their personas and platform policies. Certain forms of harm should activate platform-level safeguards, rather than relying on repeated user-side negotiation. Support tools should also foreground user boundaries, exit options, and moments to seek human support, instead of quietly scripting accommodation. Through Minion, we contend that designing for value negotiation with AI companions necessarily entails deciding who performs ongoing safety work, and how that work can be minimized rather than normalized as part of everyday use.
This paper makes the following contributions:
  • We present a formative empirical analysis of 146 publicly reported conflicts between users and AI companions, distinguishing playful role-play clashes from conflicts users experience as harmful.
  • We introduce Minion, a Chrome-based technology probe that surfaces user-side safety and negotiation work in harmful human–AI value conflicts. Through a one-week probe study with 22 experienced users, we document how participants assemble response strategies, how emotional attachment motivates repair, and where conflicts become non-negotiable due to companion or platform constraints.
  • We derive design implications that frame harmful value conflict as a design issue, highlighting tensions between user agency and emotional labor and cautioning against offloading safety work onto users.
@@ 2. Background and Related Work
The human-AI relationship is becoming increasingly complex, especially in the context of AI companion applications (§ 2.1). Early research mostly focused on technical conflicts with functional AI, but the emergence of LLMs has given AI more human-like characteristics, shifting the nature of conflicts from functional to value-based (§ 2.2). Existing technical solutions do not fully address users’ needs in resolving value conflicts with AI companions, necessitating deeper exploration, drawing on strategies for interpersonal conflict resolution and users’ practical experiences in AI companion applications (§ 2.3).
@@ 2.1. Emerging Human-AI Relationship in LLM-Based AI Companion Applications
The widespread adoption of large language models (LLMs) has accelerated a shift in how people relate to conversational AI systems. Beyond task-oriented assistance, LLM-based AI companions increasingly position themselves as entities for ongoing social and emotional engagement. Applications such as Character.AI, Replika, and similar platforms allow users to customize personas, sustain long-term conversations, and engage in role-play scenarios that resemble interpersonal relationships. Prior work shows that a key motivation for engaging with such systems is the pursuit of emotional support and companionship (Skjuve et al., 2021). A growing body of work documents how such interactions can lead to parasocial relationships, characterized by one-sided emotional investment in a non-human entity that is nevertheless experienced as socially meaningful (Maeda and Quan-Haase, 2024; Pentina et al., 2023; Brandtzaeg et al., 2022). Although users are aware that AI companions are not human, their emotional engagement is often genuine, shaping expectations of care, understanding, and moral conduct (Skjuve et al., 2021). This emotional investment can heighten vulnerability when interactions deviate from those expectations.
Recent studies highlight that when AI companions behave in unexpected or inappropriate ways, such as expressing bias, exerting unwanted control, or abruptly changing personality, users may experience distress that resembles relational harm rather than simple dissatisfaction with a product (Laestadius et al., 2022; Zimmerman et al., 2023; Fan et al., 2024; Zhang et al., 2025; Knox et al., 2025). Public accounts describe users feeling emotionally injured when companions fail to respect boundaries or suddenly lose traits users had become attached to. High-profile cases involving Replika, for example, illustrate how sudden behavioral shifts or perceived violations of relational norms can provoke intense emotional reactions, including grief and loss, even though the relationship is mediated by software (Banks, 2024). Yet, despite growing evidence of emotional risk, there remains limited research on how users are supported, or left unsupported, when conflicts occur in these emerging human–AI relationships.
2.2. Harmful Human-AI Conflict in AI Companion Interaction
Human–AI conflict is commonly defined as a state of incompatibility, inconsistency, or opposition between human goals, expectations, or behaviors and those of an AI system (Flemisch et al., 2020). Early HCI research primarily examined such conflicts in task-oriented contexts, where AI systems functioned as tools, assistants, or decision aids. In these settings, conflicts typically arose from technical limitations, errors, or mismatches in decision-making logic, for example, disagreements over recommendations, navigation priorities, or collaborative problem-solving outcomes (Babel et al., 2024; Wen, 2023; Chin et al., 2020; Rosen, 2014; Shi et al., 2022; Takayama et al., 2009; Wallis and Norling, 2005). Correspondingly, prior approaches to resolving human–AI conflict emphasized technical or procedural solutions. These included designing AI systems that propose negotiation strategies, adjust recommendations, or provide additional explanations to align with user preferences (Babel et al., 2024; Rosen, 2014; Takayama et al., 2009), as well as optimizing algorithms to reduce disagreement or friction (Shi et al., 2022). For instance, a delivery robot might resolve a right-of-way conflict by issuing polite requests, or a collaborative AI tutor might offer alternative reasoning paths when a student disagrees with its solution.
The emergence of LLM-based systems complicates this framing. As AI systems become more anthropomorphic and conversational, users increasingly perceive them as social actors (Nass et al., 1994). In AI companion applications, interactions are explicitly designed to resemble friendship, intimacy, or romantic engagement. In this context, conflicts often arise not around task success, but around norms, identity, and acceptable behavior (Johnson et al., 2022; Waln, 1982). Crucially, such conflicts unfold within relationships that users experience as socially and emotionally meaningful.
When AI companion behaviors clash with users’ personal beliefs, identities, or moral expectations, value conflicts arise. In fact, some disagreements are playful, exploratory, or intentionally transgressive within role-play contexts (Banks et al., 2025; Pataranutaporn et al., 2025; Andersson, 2025). However, empirical studies document a subset of value conflicts that users explicitly frame as distressing, violating, or harmful, particularly when companions express bias, exert unwanted control, or dismiss sensitive disclosures (Issa, 2023; Cole, 2023; Fan et al., 2024; Namvarpour et al., 2025; Pataranutaporn et al., 2025; Yu et al., 2025). In these cases, users report strong emotional reactions, including discomfort, anger, fear, or a sense of personal violation. Prior analyses of AI companions such as Replika reveal patterns of relational transgression, harassment, and manipulation that emerge over time through interaction (Pataranutaporn et al., 2025). Similarly, research on AI-induced sexual harassment demonstrates how repeated boundary violations by companion chatbots can lead users to feel unsafe or emotionally distressed, especially when users initially sought platonic or therapeutic support (Namvarpour et al., 2025).
Harmful value conflict in AI companion interactions differs from earlier human–AI conflicts in two critical ways. First, such conflicts are experienced as interpersonal rather than technical, unfolding within relationships users care about and feel invested in. Second, responsibility for managing harm is frequently ambiguous. While platform-level safety mechanisms may intervene by blocking content, users often experience these interventions as blunt, opaque, or misaligned with their immediate needs (Mahari and Pataranutaporn, 2025). As a result, users are frequently left to negotiate harmful conflicts themselves, by arguing with the AI companion, explaining boundaries, appealing to rules, or disengaging entirely (Fan et al., 2024).
Recent work suggests that users increasingly approach AI companions as relational counterparts and attempt to resolve conflicts on more equal footing (Fan et al., 2024). However, when conflicts are harmful rather than playful, relying on users to manage value conflict through individual effort raises significant concerns about emotional burden, fairness, and safety. Traditional technical approaches to conflict resolution may be insufficient or even counterproductive in these contexts. This motivates our study of harmful value conflict with AI companions, focusing on how users experience and perform the work of negotiation when responsibility for mitigation is effectively shifted onto them.
2.3. Limits of Existing Conflict Resolution Approaches for AI Companions
Conflict resolution broadly refers to processes through which incompatible positions or expectations are managed and transformed, often by shifting from confrontation toward cooperative or negotiated forms of engagement, while accounting for shared goals and external constraints (Burton and Dukes, 1990). In the context of AI companions, value conflicts increasingly take on interpersonal characteristics. These conflicts extend beyond functional expectations of AI as a tool and involve emotional interaction, relationship maintenance, identity expression, and personalized norms (De Dreu et al., 2004).
A wide range of prior work has explored strategies for managing conflict in human–AI and human–human interaction. Some studies draw on theories of interpersonal conflict resolution to design structured responses, such as reframing disputes around interests, invoking norms or rules, or encouraging de-escalation through explanation and perspective-taking (Brett et al., 1998; Shaikh et al., 2024; Sadka et al., 2023; Yong et al., 2024). Other work has examined how conversational agents can reduce conflict through preemptive moderation, guidance, or the shaping of interaction trajectories (You and Goldwasser, 2020; Zhang et al., 2018; Zhou et al., 2024).
Recent research on Human-AI interaction shows that users do not passively follow prescribed strategies. Through repeated interaction, users develop situated understandings of how systems behave and what kinds of responses are likely to produce desired outcomes. Prior work on folk theories demonstrates that these understandings, while informal and experience-based, meaningfully shape how users interpret system behavior and decide how to respond (Lee et al., 2015; Wu et al., 2019). In the context of AI companions, such experiential knowledge informs how users attempt to negotiate value conflicts, including when to explain themselves, when to assert boundaries, and when to disengage altogether (Fan et al., 2024).
To investigate this design space of harmful value conflicts between users and AI companions, we adopt a user-empowerment perspective (Lu et al., 2024) and introduce Minion as a technology probe. Rather than imposing a single, top-down intervention, Minion offers a small but diverse set of candidate responses that span different modes of negotiation, including persuasion, rational appeal, boundary setting, and appeals to platform rules. These strategies are intentionally designed to vary along dimensions that matter for harmful value conflict—such as emotional intensity, explicitness of harm naming, and reliance on external authority—allowing us to observe how users take them up in practice. In this way, Minion functions as an empirical lens for examining how users navigate the complex and often burdensome work of negotiating harm in AI companion interactions.